A two digit number is 4 times to its sum of digits , when 9 is added to the number, the digits will get reversed. Then what is that number?

Grade:

1 Answers

Sudheesh Singanamalla
13 years ago

let 2 digit number is 10x+y where x is tens and y is ones place.

4*(x+y) =  10x + y

==> 4x + 4y = 10x + 10y

==> -6y = 6x

x + y  = 0 ------------------------ (1)

10x + y + 9 = 10y + x

9x - 9y = -9

x-y = -1 ----------------------------- (2)

solving (1) & (2) we get

2x = -1

x = -1/2

y = 1/2

number = 10y+x = 5 - 1/2 = 9/2